OpenStudy (anonymous):

please help: |-3(3) -4| -1

OpenStudy (anonymous):

Start by distributing the -3 and subtracting 4. What do you get?

OpenStudy (anonymous):

\[\left| -3\left( 3 \right)-4 \right|-1=\left| -9-4 \right|-1=13-1=12\]

OpenStudy (anonymous):

I see what I did there. 12 is the correct answer.